OpenStudy (anonymous):

2/3 x +16= -14 What is the solution to the equation?

OpenStudy (anonymous):

First you subtract 16 from both sides Then you get rid of the 2/3 by multiplying by the reciprocal (3/2)

OpenStudy (anonymous):

or multiply the entire equation by 3, to get rid of the fraction, and then solve for x

OpenStudy (anonymous):

can't do that....might get suspended

OpenStudy (anonymous):

people get suspended all the time too

OpenStudy (anonymous):

just this once... 2/3x + 16 = -14 (multiply the entire equation by 3) 2x + 28 = - 42 (subtract 42 from both sides) 2x = -42 - 28 2x = - 70 x = -35

OpenStudy (anonymous):

thank you...

OpenStudy (anonymous):

wait...I messed up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x + 48 = - 42 (subtract 48 from both sides) 2x = -42 - 48 2x = - 90 x = -45 check... 2/3(-45) + 16 = -14 -90/3 + 16 = -14 -30 + 16 = -14 -14 = -14 (correct) x = -45

OpenStudy (anonymous):

if your wrong...

OpenStudy (anonymous):

I am not...x = -45

OpenStudy (anonymous):

Okay you where right

OpenStudy (anonymous):

:)